International coffee Grounds has two kinds of coffee. Coffee a costs $12 per kilogram, and Coffee B costs $8 per Kilogram. How many kilograms of each should be combinedto get 100 kilograms of a blend worth $11 per kilogram

Answer by jorel1380(3719)

A+B=100
12A+8B=1100
8A+8B=800
4A=300
A=75
B=25
You need 75 kgs. of coffee A and 25 kgs. of coffee B..